22. Go Dawgs makes a product with the following standard costs lbs or hrsStnd Price/Rate needed Per Unit per Unit of Measure Direct Materials (lbs) Direct Labor (hrs) Variable Overhead Costs 21.00 In June the company peoduced 8,500 units using 36,125 required 12,750 direct labor-hours to build those units. The actual direct labor rate was $22.00 per hour and the lotal actual variable overhead costs were $43,350. The materials price variance is compuned when maserials are used. Variable overhead is applied on the basis of direct labor-hours the company peoduced 8,500 units using 36,125 pounds at a total direct material cost of $126438 and Required a. Compute the product cost variance, in total. b. Compute the materials quantity variance e. Compute the materials price variance d. Compute the labor efficiency variance e. Compute the labor rate variance. f Compute the variable overhead efficiency variance g Compute the variable overhead rate variance 23. Marchmen Corp has provided the following information Selling Price per unit of S120, Variable Expense per unit of $72, sales quantity of 4,000 units and fixed expenses or $166,000 Answer the following questions: (refer back to the base case scenario for each question, below) a. What is the break-even number of units? b. Management is considering using a new component that would increase the unit variable cost by $2 Since the new component would increase the features of the companys product, the marketing manager predicts that sales would increase by 200 units. What should be the overall effect on the companys net operating income of this chaoge? c. The marketing manager would like to introdace asies commissions as an incentive for the sales staff The marketing manager has proposed a cosion of $8 per unit. In exchange, the sales staff would accept a decrease in their salaries of $27,000. (This is the companys savings for the entire sales staff) The marketing manager predicts that introducing this sales incentive would increase sales by 100 units What should be the overall effect on the companys net operating income of this change? d. The company MUST SELL all of the 5,000 units left in inventory. What is the minimum sales price they can charge and still break even?
Answer #1

Answer to Question 22.

Product cost variance = Standard cost - Actual Cost

Standard cost Actual cost
Direct material cost $102000 ($4 × 3 × 8500 units) 126438 ( given )
Direct labor cost $357000 ($21 × 2hours × 8500 units) $280500($22 × 12750 hours )
Variable overhead cost $59500 ( $3.5 × 2 × 8500 units ) $43350 ( given )
Total $518500 $450288

= $518500 - 450288 = 68212 F.

Material Quantity Variance = Standard price × ( standard quantity - Actual quantity)

= $4 × ( 25500( 8500 × 3 ) - 36125) = $42500 A

Material Price Variance = Actual Quantity × ( Standard price - actual price )

= 36125 ( $4 - $3.5 ) = $18062.5 F.

Labor Efficiency Variance = standard rate × ( standard hours - actual hours )

= $21 × ( 17000 - 12750 ) = $89250 F.

Labor Rate Variance = Actual time × ( standard rate × actual rate )

= 12750 × ( $21 - $22 ) = 12750 A.

Variable Overhead Efficiency Variance

= Standard rate × ( standard hours - actual hours )

= 3.5 × ( 17000 - 12750 ) = $14875 F.

Variable Overhead Rate Variance

= Actual hours × ( standard rate - actual rate )

= 12750 × ( $3.50 - $3.40( 43350 ÷ 12750))

= 12750 × $0.10 = $1275 F.